OpenStudy (anonymous):

Solve the system by the method of substitution. Solve system by graphing 3y = –1/2x + 2 y = –x + 9

OpenStudy (anonymous):

Will give medal

OpenStudy (goldeverything):

for this question do you need a graph

OpenStudy (anonymous):

no

OpenStudy (goldeverything):

ok so you just need help with the answer

OpenStudy (anonymous):

yes plz

OpenStudy (goldeverything):

what math subject are you doing at the moment

OpenStudy (anonymous):

Algebra 2

OpenStudy (goldeverything):

ok now i know what to do

OpenStudy (goldeverything):

just before i start are there option choices

OpenStudy (goldeverything):

ok nvm so you start of by combining the equation

OpenStudy (goldeverything):

\[4y = -1/2x+12\]

OpenStudy (anonymous):

o sry u left so i went to grab a bite and yes there are ansew choices

OpenStudy (anonymous):

*answer

OpenStudy (goldeverything):

no worries its fine

OpenStudy (anonymous):

what how did u get the 4y equation

OpenStudy (goldeverything):

combine 3y and y, y by itself equals 1

OpenStudy (anonymous):

(3, 6) (20, –4) (10, –1) (–1, 8)

OpenStudy (anonymous):

those are the answer choices

OpenStudy (goldeverything):

so nvm don't combine anything

OpenStudy (goldeverything):

solve it one step by a time

OpenStudy (anonymous):

plug them into the equation

OpenStudy (goldeverything):

\[3y =1/2x +2\]

OpenStudy (goldeverything):

try solving that

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

i cant because 1/2 divided by 3

OpenStudy (blurbendy):

you know y = -x + 9, so substitute that in for y in the first equation

OpenStudy (anonymous):

substitute what for y

OpenStudy (blurbendy):

the bottom equation tells you y = -x + 9 so put (-x + 9) in wherever you see y in the top equation

OpenStudy (anonymous):

o kk

OpenStudy (blurbendy):

what do you get?

OpenStudy (anonymous):

-3x-27=-1/2x+2

OpenStudy (blurbendy):

3(-x + 9) = -3x + 27 = -1/2x + 2 you were close, just flipped the signs in front of the 27

OpenStudy (blurbendy):

so no you can solve for x

OpenStudy (blurbendy):

now*

OpenStudy (anonymous):

so in order for me to solve for x just take the point and put the in

OpenStudy (anonymous):

sry im still a little confused

OpenStudy (blurbendy):

so now you have -3x + 27 = -1/2x + 2 and you want to solve for x now so first get the x's on the left side, and move everything else to the right side

OpenStudy (anonymous):

so i would add 1/2

OpenStudy (blurbendy):

yup, do what's -3 + 1/2 ?

OpenStudy (blurbendy):

so*

OpenStudy (anonymous):

-2 1/2

OpenStudy (blurbendy):

yes, which is -5/2, so now you have -5/2x + 27 = 2 you have to get x by itself so subtract 27

OpenStudy (anonymous):

- 27 5/2 = 2 ?

OpenStudy (blurbendy):

-5/2x = -27 + 2

OpenStudy (anonymous):

lol sry

OpenStudy (blurbendy):

it's okay. what does the equation look like now?

OpenStudy (anonymous):

i put it in decimal form bcz i suck at fractions and i got 29.5

OpenStudy (blurbendy):

-27 + 2 = -25 so you have -5/2x = -25

OpenStudy (anonymous):

so for me to get x by itself just add 5/2

OpenStudy (blurbendy):

no, you first have to multiply both sides by 2 so you get -5x = -25 * 2

OpenStudy (anonymous):

? sry im confused you don't have to keep helping if u dont want to sry

OpenStudy (blurbendy):

when you have a fraction in front of a variable, like -5/2x = -25 you need to get x by itself right? so you have first mulltiply both sides by 2 to get ride of the 2 on the bottom so youre left with -5x = -25 * 2 = -50

OpenStudy (blurbendy):

so then you have -5x = -50 but you still have a -5 in front, so you have to divide by -5 now x = -50 / -5 x = 10

OpenStudy (anonymous):

Thank you sry for taking so much of your time

OpenStudy (blurbendy):

it's okay. youre not done yet though. you have to put x = 10 in to x in the bottom equation and solve for y.

OpenStudy (anonymous):

o kk

OpenStudy (blurbendy):

what did you get?

OpenStudy (anonymous):

19

OpenStudy (blurbendy):

y = -(10) + 9 y = -10 + 9 y = -1

OpenStudy (anonymous):

o the negative

OpenStudy (anonymous):

thx

OpenStudy (blurbendy):

yeah, gotta be careful with those

OpenStudy (anonymous):

lol

OpenStudy (blurbendy):

you're welcome!